Question: A firm currently employs 40 production workers and 5 supervisors. The marginal product of the last production worker employed is 60 units of output per hour and production workers are paid $8 per hour. The marginal product of the last supervisor employed is 120 units of output per hour and supervisors are paid $20 per hour. Every employee works 40 hours per week.

a) What is the firm's total labor cost per week?

b) Assume that the firm's isoquants are smooth curves and that labor hours can be varied continuously. Is the firm producing the maximum level of output given its current level of cost? Explain why or why not. If it isn't, explain what it should do to increase output.
Answer #1

A)40 workers and 5supervisors

Total labor cost=40*8+5*20=320+100=420 for 1 hour

Thus every employer works for 40hours=420*40=16800

B)For cost minimisation

marginal product of last dollar from each factor should be same

thus 60/8=7.5>120/20=6

Thus firm is not minimising cost

Thus firm should hire more worker and less supervisors in order to minimise cost.